A circle placed between outer square and inner square. the length of the outer square wall is 10. How to find formula of the circle radius and length of inner square wall?
Answer by Theo(13342)

the diameter of the circle will be equal to the length of the outer square wall.

the diameter of the circle will be equal to the length of a diagonal of the inner square.

the length of the outer square wall is equal to 10.

the diameter of the circle is also equal to 10.


the length of the inner square wall is equal to square root of 50.

this is because the diameter of the smaller square is a hypotenuse of the right triangle that has the side of the smaller square as its legs.

if we allow the length of the inner square wall be equal to x, by pythagorus, we get x^2 + x^2 = 10^2 which becomes 2x^2 = 100 which becomes x^3 = 50 which becomes square root of x = square root of 50.

this assumes the inner square is inscribed in the circle and the outer square has the circle inscribed into it.

the following diagram shows you what i believe is the problem you posted.

$$$